OpenStudy (anonymous):

The function f(t)=5/3e^(-5t/3) models a continuous probability for 0

OpenStudy (anonymous):

\[0 \le t < \infty \]

OpenStudy (anonymous):

Notice that \[ \int_0^x \frac{5}{3} e^{-\frac{5 t}{3}} \, dt=1-e^{-\frac{5 x}{3}} \] Find x so that \[ 1-e^{-\frac{5 x}{3}}=\frac 1 2 \] and you are done

OpenStudy (anonymous):

\[ 1-e^{-\frac{5 x}{3}}=\frac 1 2\\ e^{-\frac{5 x}{3}}=\frac 1 2\\ -\frac{5 x}{3}=-\ln (2)\\ x=\frac{3 \ln(2) } {5} \]

OpenStudy (anonymous):

Interesting, the median for a discrete function and for a continuous function don't seem to correlate much at all.